### v1

# Bohao Wu   Qingyun Wang   Yue Guo

Abstract
Personalizing jargon detection and explanation is essential for making technical documents accessible to readers with diverse disciplinary backgrounds. However, tailoring models to individual users typically requires substantial annotation efforts and computational resources due to user-specific finetuning. To address this, we present a systematic study of personalized jargon detection, focusing on methods that are both efficient and scalable for real-world deployment. We explore two personalization strategies: (1) lightweight fine-tuning using Low-Rank Adaptation (LoRA) on open-source models, and (2) personalized prompting, which tailors model behavior at inference time without retaining. To reflect realistic constraints, we also investigate hybrid approaches that combine limited annotated data with unsupervised user background signals. Our personalized LoRA model outperforms GPT-4 by 21.4% in F1 score and exceeds the best performing oracle baseline by 8.3%. Remarkably, our method achieves comparable performance using only 10% of the annotated training data, demonstrating its practicality for resource-constrained settings. Our study offers the first work to systematically explore efficient, low-resource personalization of jargon detection using open-source language models, offering a practical path toward scalable, user-adaptive NLP system.
Explain Less, Understand More: Jargon Detection via Personalized Parameter-Efficient Fine-tuning
Bohao Wu   Qingyun Wang   Yue Guo University of Illinois at Urbana-Champaign {bohaowu,qingyun4,yueg}@illinois.edu
1 Introduction
Figure 1: Bridging the disciplinary gap under constraints of limited compute and annotation: Our approach leverages PEFT, prompting, and mixed data to enable efficient and effective adaption.
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide spectrum of applications, showcasing sophisticated natural language understanding and generation ability Zhou et al. (2024); Wang et al. (2024); Baek et al. (2025). As research increasingly transcends disciplinary boundaries, LLMs are being leveraged to facilitate the integration of diverse fields, offering new avenues for processing complex, cross-disciplinary data and generating novel insights Leto et al. (2024); Ramoneda et al. (2024); Lu et al. (2024); Jiang et al. (2025).
However, a significant impediment to seamless interdisciplinary collaboration and knowledge sharing lies in the challenge of domain-specific jargon Barnett and Doubleday (2020). Researchers often encounter specialized terminology outside their core expertise, which can hinder effective communication and understanding across disciplines Han et al. (2018); Lucy et al. (2023). This is particularly salient when scholars need to engage with literature from unfamiliar domains, potentially leading to misinterpretations and impeding the synthesis of knowledgeGlasziou et al. (2020). Our primary motivation is to provide personalized assistance to researchers in navigating this linguistic landscape, enabling a more accessible and efficient engagement with interdisciplinary content. A crucial first step in realizing this goal is the accurate identification of domain-specific jargon tailored to the individual researcher’s background and expertise. Therefore, this paper focuses on personalized jargon identification.
Figure 2: Schematics of the supervised, unsupervised and the combined training strategies for personalized LLMs.
Additionally, the increasing demand for user-centric AI has spurred significant interest in personalized Large Language Models (LLMs). While various strategies have been proposed, each presents its own limitations. Benchmarks like LaMP (Language Model Personalization) (Salemi et al., 2024) underscore the importance of this area but also reveal the complexities in achieving effective personalization across diverse tasks. Techniques focusing on adapting the model itself, such as One PEFT Per User (OPPU) (Tan et al., 2024b) and Personalized Pieces (Per-Pcs) (Tan et al., 2024a), demonstrate promising results in encoding user preferences through parameter-efficient fine-tuning or lightweight plug-in modules. However, these methods typically rely on supervised data derived from explicit user interactions or preferences, which can be sparse, costly to obtain, and may not fully generalize to unseen scenarios or capture implicit user needs. Furthermore, while efforts like HLLM (Chen et al., 2024) explore personalization in specific downstream tasks like recommendation, the applicability of such task-specific architectures to the broader challenge of personalized language understanding remains an open question. The need for data-efficient and broadly applicable personalization, especially when explicit user feedback is scarce, motivates our exploration of unsupervised personalized PEFT for jargon identification.
Addressing the challenges at the intersection of interdisciplinary communication and personalized language models, Guo et al. (2024) pioneered the use of LLMs, employing GPT-4 OpenAI (2023) to gauge out-of-domain scholars’ familiarity with jargon. Building upon their innovative work, which faced computational costs and suboptimal performance, our research aims to overcome these limitations by developing more efficient and tailored solutions leveraging lightweight, open-source LLMs for individual researchers. Furthermore, recognizing the data-intensive nature and potential generalization limits of prevalent supervised personalized PEFT approaches, this paper investigates the potential of unsupervised personalized PEFT to enhance jargon identification. Our contributions are:
  • Personalized Language Model Finetuning: We introduce a PEFT approach with LoRA that significantly improves personalized jargon identification over prior work (Guo et al., 2024) with minimal resources.
  • Personalized Prompting Strategies: We demonstrate that tailored prompting based on researcher background significantly influences jargon identification across domains, offering efficient performance optimization.
  • Personalization through Unsupervised Background Integration: We use unsupervised researcher data with minimal labeled data (10%) to achieve performance comparable to fully annotated models.
2 Method
2.1 Prompting Setup
We retrieve a series of background contexts to guide our personalized jargon identification. Specifically, given a pair of the research abstract and the corresponding entity of interest extracted from it, we retrieve four types of background contextual information (details in App. Table 2): 1) Metadata: We retrieve the annotator’s metadata from Semantic Scholar API Lo et al. (2020), including the abstract’s domain, the annotator’s publication number, the annotator’s reference count, and the annotator’s definition of their research domain. 2) Profile: As illustrated in Fig 2, we synthesize the researcher profile using the baseline model, which describes the annotator’s background and research information based on his/her publications. 3) Nearest Annotator: We leverage BM25 (Trotman et al., 2014) as our retrieval function, with the collection of annotator profiles serving as our document library, which we term the profile network. This method retrieves the annotator most similar to each target within the profile network, using their familiarity labels as contextual information for the target. 4) Nearest Abstract: Similarly, we apply BM25 to all the abstracts in the dataset, which we refer to as the abstract network. Then, we retrieve the most similar abstract within this network and use the target annotator’s familiarity labels as the contextual information. 5) For comparison, we also add a vanilla prompting, which means there is no input for related data, so the model has no information about the annotator.
2.2 Supervised Model Training
Based on the findings of Tan et al. (2024b), which demonstrate that LoRA (Hu et al., 2022) consistently outperforms other parameter-efficient fine-tuning (PEFT) methods, we adopt LoRA for model adaptation. Our objective in this stage is to assess whether applying LoRA to the complete dataset can lead to performance improvements beyond those achieved in previous studies (Guo et al., 2024). In this supervised setting, we maintained a consistent prompting strategy for both training and inference, aligning with the specific design of each context enhancement method detailed in Table 3. Following the Alpaca template (Taori et al., 2023), each input includes a jargon entity along with its corresponding abstract, placed in the “Input”, while the binary familiarity label is placed in the “Response”.
2.3 Unsupervised + Personalized PEFT
To assess the performance of unsupervised PEFT in our task, we deliberately avoid using task-specific formats during fine-tuning to prevent contamination. Instead, we compile a dataset composed of the titles and abstracts from each annotator’s previously published papers. We then apply LoRA fine-tuning for a few epochs. In cases where an annotator lacks sufficient publication data for effective training, we supplement the dataset by randomly retrieving papers from their self-defined research subdomain. This setup simulates a real-world scenario in which a new researcher has little or no labeled data available. Our goal is to explore whether a personalized model can still be trained effectively under such constraints, using only domain-relevant but unlabeled content.
3 Experiments and Results
3.1 Supervised Finetuning and Familiarity Classification
Figure 3: Model performance in familiarity classification for supervised versus unsupervised models, with various fine-tuning epochs under different prompting strategies. Evaluation is done upon the validation dataset. Each node is a combination of $1$ personalized models, specifically finetuned on 11 corresponding annotators. Here, GPT-4 and Oracle demonstrate previous SOTA model performance, denoting prompt-based GPT-4 and Oracle classifiers.
In Fig. 3 (left), we present the results of fine-tuning our baseline model, Llama-3.1 8B Instruct 4bit Dubey et al. (2024), using the full annotation dataset over various training epochs. Across all prompting strategies, a rapid performance increase is observed within the initial 20 epochs, followed by a plateau. Vanilla prompting achieves the highest effective F1 score of 0.82. Metadata-enhanced and profile-enhanced prompting exhibit similar performance, saturating around 0.80. Conversely, the nearest annotator and nearest abstract prompting strategies show lower performance, leveling off at approximately 0.75. All evaluations were conducted on the validation dataset. For comparison, Figure 3 (left) shows that our LoRA-based models significantly outperform prior state-of-the-art results from Guo et al. (2024)) (GPT-4 context learning: F1 0.64; oracle classifier: F1 0.72), even with simpler prompting strategies, demonstrating the effectiveness of supervised fine-tuning with LoRA in learning semantic connections between jargon and abstracts.
3.2 Unsupervised Finetuning
The performance of our unsupervised fine-tuned models, evaluated on the validation dataset across varying epochs and prompting strategies, is illustrated in Fig. 3 (right). The results indicate an initial performance dip upon commencement of fine-tuning, followed by a subsequent improvement, reaching its peak at 50 training epochs. Notably, the nearest annotator and nearest abstract prompting strategies yielded nearly 100% mismatch rates due to the unsupervised models’ lack of explicit task format knowledge and the longer context lengths associated with these strategies. Therefore, their results are excluded from the figure. Overall, unsupervised fine-tuning demonstrates a modest improvement over the baseline pre-trained models but does not achieve the performance levels of state-of-the-art supervised methods.
Figure 4: Validation performance plots for models fine-tuned under different strategies. To enable a consistent comparison across strategies, each involving datasets of varying text lengths, we use training steps as the x-axis. This also provides an approximate indication of the computational cost associated with each fine-tuning approach.
3.3 Mixed Finetuning on Supervised and Unsupervised Data
This section investigates the extent to which our fine-tuned models learn semantic knowledge versus simply aligning to the task template and explores the utility of unsupervised data when labeled data is scarce. To assess semantic learning, we observed a significant decline in performance saturation when training data was reduced to 1% and 10% (Fig. 4), with 1% models underperforming unsupervised methods. This performance decline, especially the 1% model underperforming unsupervised methods, indicates that the supervised models have learnt from the semantic information within the annotated content, rather than merely aligning to the task template. To address the challenge of limited annotation, we explored a mixed strategy, finding that integrating unsupervised researcher background data with just 10% of the original training set achieved a comparable performance (up to 79.1%) to fully supervised models, significantly outperforming models trained solely on the 10% subset. This demonstrates that leveraging unsupervised data with minimal annotation can yield competitive results, reducing the burden of extensive manual labeling.
4 Conclusions
To address the barrier of domain-specific jargon in interdisciplinary communication, we presented a novel framework for personalized jargon identification. By adapting lightweight LLMs with LoRA, we significantly outperformed prior work with minimal resources. We also demonstrated the substantial impact of personalized prompting based on the researcher’s background. Specifically, our data-efficient technique leveraging unsupervised data achieved near fully supervised performance with minimal annotation. This work offers a resource-conscious and scalable approach to building personalized tools that lower barriers to interdisciplinary understanding.
Limitations
One limitation of our current work is the reliance on a specific dataset focused primarily on computer science researchers and a limited number of out-of-domain areas. While this allowed for a controlled evaluation of our personalized techniques, the generalizability of our findings to a broader range of interdisciplinary domains and diverse researcher backgrounds requires further investigation. Future work should explore the application and evaluation of our framework on more heterogeneous datasets encompassing a wider spectrum of academic disciplines and research profiles to assess its robustness and adaptability in more varied real-world scenarios.
Ethical Considerations
In this paper, we utilized anonymized data from a pre-existing dataset, raising ethical considerations regarding the privacy and responsible use of researcher background information in future implementations. We acknowledge the potential for our personalized models to inherit or amplify biases present in pre-trained models or training data, necessitating careful evaluation across diverse user groups to ensure equitable performance. Furthermore, we recognize the importance of clarifying jargon without oversimplification and the potential for over-reliance on such tools to impact researchers’ own interdisciplinary language development. Finally, we advocate for responsible development to prevent unintended consequences like the creation of echo chambers. Ongoing evaluation and community discussion are essential to navigate these ethical complexities.
Appendix A Setup
Dataset
We use data from  Guo et al. (2024), which includes four personalized classification tasks: term familiarity and the need for additional information in the form of a definition, background, or example. Each is framed as a binary classification problem. The dataset contains 1,100 annotations per task, collected from 11 researchers who each labeled 100 out-of-domain abstracts, selecting 10 terms per abstract. The dataset also includes researcher metadata, publication records, and related background information; further details can be found in the original paper. We follow the original 60/20/20 split for training, validation, and testing to ensure direct comparability with their reported baselines.
Evaluation Metrics
To evaluate the performance of our personalized jargon identification models, we focus on the task of predicting binary familiarity labels (0 for familiar, 1 for unfamiliar) for entities extracted from research paper abstracts. Our primary evaluation metric is the F-1 score. However, During our initial baseline model selection phase, we observed that some models struggled to consistently produce the required binary label lists without additional text or nonsensical information. To account for this, we introduced the Effective F-1 Score. This metric incorporates the "Mismatch Rate" – the proportion of model outputs that did not conform to the expected binary label format. The Effective F-1 Score is calculated as follows:
[…]
Baseline Model Selection
Figure 6: Evaluation results for model selection. Here the inference is done with listed entities to ensure whether the model understands the question.The evaluations are based on the entire dataset.
To establish a robust foundation for our personalized jargon identification task using Parameter Efficient Fine-Tuning (PEFT), we first selected a suitable open-source Large Language Model (LLM) as our baseline.
We evaluated several state-of-the-art options, including Llama-2 (Touvron et al., 2023), Llama-3 (Dubey et al., 2024), Mistral v0.3 (Jiang et al., 2023), and Qwen-2.5 (Yang et al., 2024; Team, 2024), considering both their base and instruction-tuned versions in 4-bit and full precision.
The results of this initial evaluation (depicted in Figure 6) revealed significant performance variations in terms of both F-1 score and mismatch rate. While Qwen-2.5 7B Instruct 4bit achieved the highest Effective F-1 Score (0.54), and Llama-3.1 8B 4bit exhibited the lowest mismatch rate (0.5%), we ultimately selected Llama-3.1 8B Instruct 4bit as our baseline for subsequent fine-tuning experiments. This decision was based on its robust performance (Effective F-1 Score of 0.49) and its demonstrated ability to follow instructions with minimal mismatches, suggesting a strong potential for effective adaptation through PEFT for our personalized jargon identification task.
Appendix B Analysis
Does the familiarity model generalize over other personalized tasks?
In this part of the experiment, we evaluate whether the finetuned models, trained on familiarity annotations, can generalize to related but unseen tasks. Specifically, we test whether the models can predict annotators’ need for additional information (e.g., definitions, background, or examples), a task structurally different from the original familiarity labeling. This setup allows us to examine whether the models have truly internalized the annotators’ knowledge levels, or if their performance is simply a result of alignment with the annotation distribution.
Figure 5 demonstrates the strong generalization of our fine-tuned models, achieving performance on definition and background knowledge tasks comparable to prior best Lasso regression models (without explicit fine-tuning on this data) and significantly outperforming them on predicting the need for additional examples. These results suggest that supervised LoRA fine-tuning effectively captures not just annotation patterns but also a robust semantic understanding of the annotators’ domain expertise.
Model Improvement Analysis in Terms of Individual Annotators
Figure 7: Personalized model Performance for individual annotators.
Taking annotator #4 as the object, a qualitative analysis of the missed and falsely detected jargon reveals several interesting patterns. The baseline model (trained with 1% training set), while showing some capability in jargon detection, struggled with terms that exhibit a combination of characteristics. Firstly, it frequently failed to identify the terms as jargons that are relatively short and composed of common words but carry highly specific meanings within a particular domain. Examples include ’Radial curves’ (Materials Science), ’Op-amp’ (Physics), and ’Domains’ (Geology). These terms, due to their brevity and seemingly ordinary components, may have been harder for the baseline to differentiate from general language use. Secondly, the baseline model had difficulty with multi-word terms where the meaning is not a straightforward combination of the individual words, but rather a more nuanced concept. This is evident in its failure to identify ’Bayesian optimal mechanism’ (Economics), ’Riemannian framework’ (Materials Science), ’Bose-Einstein condensate’ (Physics), ’Psychometric properties’ (Economics), ’Dialectical quality’ (Philosophy), ’Explanatory account’ (Linguistics), ’Long-range ordered coupling’ (Physics, Materials Science), and ’Qualitative spatio-temporal inferences’ (Psychology). In these cases, the model may have lacked the ability to capture the semantic relationships and contextual dependencies necessary for accurate identification. Thirdly, the baseline also missed acronyms like ’CW-SSIM’ (Agricultural And Food Sciences), ’MANOVA’ (Education), and ’ARMAX model’ (Business, Engineering). Acronyms often present a challenge due to their condensed nature and lack of explicit semantic clues. Finally, there were instances where the jargon term spans multiple disciplines, such as ’Monolayers’ (Engineering, Biology), ’Peri-implant bone density’ (Materials Science, Medicine, Biology), and ’Regulatory mechanisms’ (Biology, Environmental Science), which might have added to the difficulty. While the improved model demonstrated a higher F1 score, indicative of better overall performance, it exhibited a tendency to produce more false positives. These false positives included terms like ’Savitzky-Golay (SG) filter’ (Environmental Science), ’Meta-analyses’ (Medicine), ’Post-test’ (Education), ’Quantitative research’ (Education), and ’Content analysis’ (Medicine). This suggests that the improved model, in its attempt to capture a broader range of jargon, may be more sensitive to terms that share some characteristics with jargon but are more commonly used or understood. This could indicate a trade-off where the improved model sacrifices some precision for increased recall, potentially overgeneralizing in certain contexts. Specifically, the improved model seems to be more prone to misclassifying statistical or methodological terms (e.g., ’Meta-analyses’, ’Post-test’, ’Quantitative research’) as jargon, possibly due to their frequent appearance in academic contexts, even when they are relatively well-understood within the broader research community.
Model Analysis in Terms of Jargon Domain
Figure 8: Supervised model performance difference for nearest abstract versus vanilla.
In this study, the best two models are selected, which are 100% TS with vanilla and nearest abstract (NAb) prompting. When comparing the two models across the ’Art’ and ’Philosophy’ domains, a nuanced performance profile emerges. In the ’Art’ domain, vanilla model exhibits a higher false positive rate, incorrectly identifying terms like ’Reactions’ and ’Stylistic’ as jargon, which NAb model correctly classifies. This suggests that vanilla model may be overly sensitive to terms that, while potentially used in art-related contexts, also have broader, common usage.
Conversely, in the ’Philosophy’ domain, NAb model faces challenges in both precision and recall. It exhibits a higher false positive rate, misclassifying terms such as ’Structural constraints’, ’Analytic philosophers’, and ’Argument Facets’. This indicates a tendency to over-identify common philosophical terms as highly specialized jargon. Furthermore, NAb model also demonstrates lower recall in the Philosophy domain, failing to detect several jargon terms, including ’Computational argumentation’, ’Corpus with 320 arguments’, ’Nonmonotonic inference methods’, ’Super-knotty rope’, ’Super-knot’, and ’Dialectical quality’. These terms represent complex philosophical concepts that NAb model struggles to recognize as domain-specific jargon.
Models F-1 Score $\uparrow$ Recall Precision
Oracle Classifier 71.9±1.7 76.0±2.1 68.2±2.1
GPT-4 64.2±1.5 98.7±0.5 47.6±1.6
1%TS 53.5±2.8 56.0±4.5 51.5±3.2
10%TS 63.6±2.9 59.8±5.1 69.0±1.7
Unsuper 54.6±5.1 77.5±7.6 45.0±0.8
Unsuper+10%TS 77.0±1.1 78.9±2.6 75.4±0.7
100%TS 77.9±1.2 77.8±2.2 78.0±0.5
100%TS+MD 76.8±1.1 76.1±2.9 77.7±1.0
100%TS+Profile 76.6±1.0 73.7±2.0 79.9±1.6
100%TS+NAn 72.1±3.0 70.3±4.1 74.5±2.7
100%TS+NAb 77.8±1.1 78.3±1.7 77.5±2.5
Table 1: Evaluation of the finetuned models on the test dataset. If not specified, the finetuning and inference are performed with vanilla prompting. Acronym: ‘TS’-traing set, ‘MD’-metadata, ‘NAn’-nearest annotator, ‘NAb’-nearest abstract.
Strategies Related data
Vanilla "" (empty string)
Metadata "Self-defined subfield of the reader is: {} Number of papers published by the reader is: {} Number of papers referenced by the reader is: {} Year of the first paper published by the reader is: {} Domain of study of the paper is: {}"
Profile-enhancement "This reader is a domain expert in natural language processing (NLP) …" (Machine-generated profile)
Nearest annotator Another researcher similar to the reader has read the same abstract. For the entity list {entity_list}, this researcher provides the familiarity list as {familiarity_list}.
Nearest abstract For another similar abstract with the entity list {entity_list}, this reader provides the familiarity list as {familiarity_list}.
Table 2: The prompting strategies for both supervised fine-tuning and inference.
Tasks Instructions Prompt
Familiarity classification Your job is to estimate how much the reader knows about an entity. You will be provided with the entity, the abstract where the entity came from, and related data about either the reader or the abstract. Your answer should be a list of binary, either 0 or 1, of the same length as the entity list, with no other words. Entity: {entity} Abstract: {abstract} Additional information: {related_data} Here’s how to gauge the reader’s familiarity: - 0: The reader knows this subject well and can describe it to others. - 1: The reader has either encountered this subject before but knows little about it, or has never come across it at all. Based on the information provided, determine familiarity score list corresponding to the entity list, a list of either 0 or 1:
Definition needs classification You are tasked with predicting whether the reader might need additional Definition/Explanation to fully grasp the entities mentioned in a given abstract. You will be provided with the entity list, the abstract where the entities come from, and related data pertinent to the reader or the abstract. Definition of definition/explanation: provides key information on the term independent of any context (e.g., a specific scientific abstract). A definition answers the question, "What is/are [term]?" Entity: {entity} Abstract: {abstract} Additional information: {related_data} Provide the estimation whether additional information is needed in a list in the order of the entity. The estimation should be either 0(no) or 1(yes). No need to mention the entity:
Background needs classification You are tasked with predicting whether the reader might need additional Background/Motivation to fully grasp the entities mentioned in a given abstract. You will be provided with the entity list, the abstract where the entities come from, and related data pertinent to the reader or the abstract. Definition of background/motivation: introduces information that is important for understanding the term in the context of the abstract. Background can provide information about how the term relates to overall problem, significance, and motivation of the abstract. Entity: {entity} Abstract: {abstract} Additional information: {related_data} Provide the estimation whether additional information is needed in a list in the order of the entity. The estimation should be either 0(no) or 1(yes). No need to mention the entity:
Example needs classification You are tasked with predicting whether the reader might need additional Example to fully grasp the entities mentioned in a given abstract. You will be provided with the entity list, the abstract where the entities come from, and related data pertinent to the reader or the abstract. Definition of example: offers specific instances that help illustrate the practical application or usage of the term within the abstract. Entity: {entity} Abstract: {abstract} Additional information: {related_data} Provide the estimation whether additional information is needed in a list in the order of the entity. The estimation should be either 0(no) or 1(yes). No need to mention the entity:
Table 3: The configuration of instructions and prompts for training and inference, following the prompting format from the previous work (Guo et al., 2024).